# On the PM2.5 – Mortality Relationship: A Bayesian Model for Spatio-Temporal Confounding

### v2

@@ Abstract
In epidemiological studies about the effects of air pollution on public health, it is common practice to incorporate an arbitrary smooth function of time (or space) in time-series (or spatial) data to account for unknown confounding effects. These smooth functions are inherently linked to exposure, thereby increasing the difficulty of assessing the exposure effect (i.e. the concurvity problem). This issue is particularly challenging in spatio-temporal analyses, where it has been less explored so far. To study the effects of fine particulate matter on mortality in Italy, a model that simultaneously accounts for spatio-temporal confounding and for the non-linear form of the effect of interest is needed. Henceforth, we propose a Bayesian spatial dynamic generalized linear model, which allows for non-linear associations and for a decomposition of the exposure effect. This decomposition accommodates associations with the outcome at fine and coarse temporal and spatial scales of variation. These features, when combined, allow reducing the spatio-temporal confounding bias and recovering the shape of the association, as demonstrated through simulation studies. The real-data analysis indicates that the exposure effect shows a clear temporal dynamic, with peaks during summer. We argue that this may be due to interactions of particulate matter concentrations with air temperature and unmeasured confounders.
Keywords: Air pollution, Bayesian, Concurvity, Dynamic generalized linear model, Non-linearity, Spatio-temporal confounding.
@@ 1 Introduction
Air pollution is a major environmental risk factor and undermines public health, causing a wide range of health problems, including premature deaths. To date, there is a very rich literature demonstrating the short- and long-term consequences of air pollution on health outcomes (see, e.g., Dominici et al. (2002), Janes et al. (2007), Zeger et al. (2008), Greven et al. (2011), Dai et al. (2014), Di et al. (2017), Wu et al. (2020), Brunekreef et al. (2021b), Renzi et al. (2021), Dominici et al. (2022), Gamerman et al. (2022), Requia et al. (2024)).
Within the ecological regression framework, one of the challenges encountered is the determination of the shape of the exposure–outcome relationship, which is commonly assumed to be constant and linear. We say that the relationship is linear if the exposure enters linearly in the specification of the link function. However, if this assumption is violated, misspecification bias would result. We acknowledge that all models are inherently flawed, but some of them are so misspecified that they can lead to incorrect inferential statements (Gelman and Shalizi, 2013).
@@
In this paper, the ultimate goal is to correctly evaluate the short-term association between particulate matter with a diameter less than $2.5\mu m$ (PM2.5) and all-cause mortality. The data span the period between 2018 and 2022 and are available for each health district of two Italian regions (see Section 2). We argue that the commonly adopted assumption of a constant and linear effect, without considering interaction with confounding variables, may be inadequate, given the spatial and temporal variability of PM2.5 and its associated health risks. Specifically: (i) the physical properties and chemical composition of the PM mixture exhibits substantial variation across locations and over time, driven by environmental conditions, human activities, and episodic phenomena such as sandstorms and wildfires (Chen et al., 2013, Kelly and Fussell, 2012, Masselot et al., 2022, Peng et al., 2005, Zhou et al., 2021); (ii) human susceptibility to PM2.5 is not static, but rather modulated by temporally and spatially varying factors, including extreme temperatures, which can exacerbate health risks in specific regions or seasons (Bergmann et al., 2020, Peng et al., 2005, Roberts, 2004, Li et al., 2017); (iii) the relationship between true and measured exposure varies across time and space, with seasonal factors like increased summer ventilation potentially reducing exposure measurement error (Bergmann et al., 2020, Peng et al., 2006, Roberts, 2004, Stafoggia et al., 2008).
To account for the potential non-linearity of the exposure–outcome relationship, various strategies are available. An exposure-response function is often estimated (see, among others, Brunekreef et al., 2021b, Cork et al., 2023, Dominici et al., 2022, Renzi et al., 2021). Alternatively, it would be appropriate to specify a model that incorporates the inherent spatial and temporal variations of the exposure effect (Gelfand et al., 2003). By doing so, we can analyze whether the strength of the relationship between PM2.5 and mortality have changed over different time periods and across health districts. The temporal dynamic of the exposure effect allows investigating the influence of various factors that have emerged since 2020, such as lockdowns, the use of facial masks, physical distancing and social restrictions. Indeed, these measures have significantly altered people’s habits and interactions with the environment. Furthermore, the health districts are characterized, for example, by morphological and climatic differences which may influence the exposure–outcome relationship. Therefore, it is important to allow for spatially varying coefficients in the model.
Time-varying coefficients models (TVCMs) have been previously considered in the literature. For example, the exposure effect is modeled either as a periodic function (Chen et al., 2013, Peng et al., 2005) or as a season-specific step function (Dai et al., 2014, Klompmaker et al., 2023). Moreover, Chiogna and Gaetan (2002) and Lee and Shaddick (2008) employ dynamic generalized linear models (DGLMs; West and Harrison, 1997) when analyzing the relationship between air pollution and mortality with time-series data.
Spatio-temporally varying coefficients models (STVCMs) are reviewed by Banerjee et al. (2014) and Sahu (2022), among others, and have been applied to various fields, including healthcare (Cai et al., 2013, Song et al., 2020), air pollution and climate studies (Paez et al., 2008, Bakar et al., 2015). This class of models provide a flexible framework to capture smooth space-time variations of the exposure–outcome association.
A second major challenge in ecological regression models relates to the problem of confounding. When data are spatially correlated, this problem has received a lot of attention in recent years — see, e.g., Reich et al. (2006), Paciorek (2010), Thaden and Kneib (2018), Dupont et al. (2022), Marques et al. (2022), Azevedo et al. (2023) and Zaccardi et al. (2025). Gilbert et al. (2023) and Khan and Berrett (2023) argue that multiple definitions of “spatial confounding” co-exist in the literature. In this paper, we extend one of their notions and refer to spatio-temporal confounding as the problematic assessment of the exposure effect due to concurvity, that is the correlation between exposure and an unknown smooth function of space-time that affects the outcome (Gilbert et al., 2023, Peng et al., 2006). In other words, factors that vary on similar spatial and temporal scales as the exposure are considered potential confounders. Although information on some of these confounders is available, such as meteorological variables, it is likely that others are unmeasured or difficult to quantify, such as influenza epidemics or changes in population structure (Peng et al., 2006). Not correctly accounting for unmeasured confounding may lead to an incorrect evaluation of the exposure–outcome relationship (i.e., confounding bias), as the model may not be able to distinguish between the effect of the exposure and that of the confounders.
Despite the complexity of this problem, just few authors have recently proposed approaches that explicitly account for spatio-temporal confounding. For instance, Reich et al. (2021) discuss extensions of causal inference methods. Other works have adopted a “data analysis” perspective (Khan and Berrett, 2023), where confounding results from the collinearity between exposure and structured random effects. Within this strand of research, Prates et al. (2022) propose the transformed multivariate Gaussian Markov random field model (TMGMRF), an extension of the approach by Prates et al. (2015), where Poisson intensities or Bernoulli rates are modeled directly and without the need of a link function. A copula structure captures the spatio-temporal dependence and guarantees the disentanglement between fixed and random effects, hence their model should be able to control for confounding bias. A spatio-temporal linear model is considered by Adin et al. (2023), where the random effects are orthogonalized with respect to the fixed effects: in other words, this work extends the restricted spatial regression model (RSR; Hodges and Reich, 2010, Reich et al., 2006). These approaches, however, are not intended to adjust for unmeasured confounders (Khan and Calder, 2022).
A comprehensive data-driven approach that can capture the temporal dynamics of the exposure effect while simultaneously controlling for spatio-temporal confounding is still lacking. To tackle both problems, we propose, under the Bayesian paradigm, the spatial DGLM for confounding (SDGLMC), which has several advantages. First, spacetime-varying intercept and slope coefficient are introduced in the model specification to account for unmeasured spatio-temporal confounders and non-linear exposure–outcome associations, respectively. Second, since DGLMs are only locally linear, they can approximate any smooth function of time (Gamerman and Lopes, 2006, Granger, 2008), hence the proposed model should not be affected by misspecification bias. Third, SDGLMC can be seen as a generalization of other methods in the literature. For example, the approaches by Chen et al. (2013) and Peng et al. (2005) cannot account for exceptional events such as the outbreak of the SARS-CoV-2 pandemic. The models proposed, among others, by Dai et al. (2014) and Klompmaker et al. (2023) assume that the effect varies between seasons, but is constant within each season. In other terms, the latter works implicitly assume the possibility of abrupt changes in the exposure effect between adjacent seasons, however these are unlikely in practice, because the exposure effect depends, e.g., on meteorological conditions and on human activity patterns that change gradually over time (Peng et al., 2005, Bhaskaran et al., 2013). In contrast, the proposed model can overcome both drawbacks, thanks to the ability of DGLMs to capture gradual changes and to emphasize both seasonal and non-systematic variations (Lütkepohl, 2005, Petris et al., 2009, Bitto and Frühwirth-Schnatter, 2019). Additionally, SDGLMC extends the approaches by Paez et al. (2008), Cai et al. (2013) and Bakar et al. (2015) since it includes a space-time interaction term in the intercept, thus enabling control for complex spatio-temporal confounding patterns.
As a fourth major advantage, following Greven et al. (2011) and Janes et al. (2007), the exposure effect is decomposed into two orthogonal components, each describing the (non-linear) association with the health outcome at a different scale of variation (see Section 3). The focus of inference is the component that captures the association at a finer scale — where confounding by unmeasured factors is less likely (Greven et al., 2011) — which is assumed to exhibit spatio-temporal dynamics. Furthermore, SDGLMC provides explicit interpretation of each part of these dynamics through an analysis-of-variance-like decomposition (see Section 3.1).
@@
This paper is organized as follows. Section 2 offers a brief description of the data that motivated this study. Sections 3 and 4 provide detailed information on the proposed model and on the inferential procedure, respectively. Section 5 presents existing approaches from the literature to which the results of the proposed model are compared. In Section 6, a simulation study is conducted to investigate the ability of SDGLMC in controlling for spatio-temporal confounding when the true association is known. The results from the real-data application are analyzed in Section 7. Finally, Section 8 provides a discussion.
2 The Data
@@
This section is dedicated to the description of the dataset that motivated our research, specifically we analyze the association between daily PM2.5 concentrations and all-cause mortality counts among the elderly (i.e. people aged 65 or more) in Piemonte and Lombardia regions (Italy) between 2018 and 2022.
@@
The data are retrieved from publicly available sources, and full details are given in Section A of the Supplementary Material. Mortality counts are collected from the Italian National Institute of Statistics (ISTAT, 2023a), and are observed on an irregular lattice comprising ${N}=117$ non-overlapping areal units termed health districts (HDs). The study period comprises ${T}=1826$ consecutive days, hence $Y_{it}$ denotes the number of deaths (by any cause) occurred in district $i=1,\dots,{N}$ during day $t=1,\dots,{T}$. The PM2.5 concentrations are used as a measure of the exposure in district $i$ during day $t$, denoted as $X_{it}$, and are collected from the Copernicus Atmosphere Monitoring Service’s (CAMS) Atmosphere Data Store (INERIS et al., 2022, CAMS, 2023). We also acquire information on temperature and relative humidity, which are potential confounders, from the ERA5 reanalyses dataset available at the Copernicus Climate Change Service’s (C3S) Climate Data Store (C3S, 2023, Hersbach et al., 2018). The remote-sensed measurements are spatially and temporally aggregated to match the resolution of the health data.
@@
A measure usually employed for exploratory analyses is the standardized mortality ratio (SMR), which is calculated as $SMR_{it}=Y_{it}/E_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$. The denominator, $E_{it}$, denotes the expected number of deaths, and is computed allowing for the underlying age–sex stratification of the population within each HD (see also Section A of the Supplementary Material). Figure 1 (left panel) depicts the average daily SMR for each HD over the years 2018–2022, whose mean value is $1.25$ over the whole study domain, and spatial clustering is noticeable. The right panel shows the long-term average PM2.5 concentrations for each HD, and on average the Lombardia region is more polluted than Piemonte ($19.41\,\mu g/m^{3}$ and $15.18\,\mu g/m^{3}$, respectively). Other descriptive statistics are reported in Section B of the Supplementary Material.
@@
Figure 1: Long-term averages over the whole study period (2018–2022) of the SMR due to all-cause mortality (left) and of the PM2.5 levels (right) for each HD. The thicker solid line separates Piemonte (the westernmost region) from Lombardia (the easternmost region).
@@
Figure 2 shows the average daily SMR (on the logarithmic scale) for Lombardia and Piemonte. The seasonal pattern is similar in both regions, with the lowest values in summer, except for the year 2020. Two peaks can be noticed, where the log(SMR) more than doubled: the first one occurred in April 2020 and the other one in November 2020. These are due to the first and second “waves” of the SARS-CoV-2 pandemic, respectively. The first wave was more severe in Lombardia, while the second one was more severe in Piemonte. In addition, the two peaks in the SMR in 2020 are not reflected in the PM2.5 time series. As can be noticed from Table B.3 in the Supplementary Material, the overall 2020 annual mean of PM2.5 levels was the lowest among the five years considered in our analysis.
@@
Figure 2: Time series of the average daily SMR (on the log scale, orange lines) and average daily PM2.5 concentrations (black lines) for Piemonte (top) and Lombardia (bottom) regions. The vertical dashed lines separate each calendar year. Blue vertical bands correspond to the national lockdown periods.
@@
Figure 2 also highlights other two aspects. One is that the highest values of the SMR in 2018 were reached at the beginning of January, suggesting that the influenza epidemic reached its peak before the end of 2017. This is different from the year 2019, where the SMR peaks were reached in February. The other aspect is that there was an increase in mortality during summer 2022. Possible reasons may be the occurrence of prolonged heat waves (indeed, the average daily temperature time-series has its absolute maximum in July 2022), or the outbreak of an influenza epidemic.
However, to study the association of interest, it is necessary to account for the effects of measured and unmeasured confounders. Failing to do so may lead to incorrect inferences, due to concurvity. The proposed approach, which is detailed in the next section, is framed to overcome this issue, among others.

### v3

# On the PM2.5 – Mortality Association: A Bayesian Model for Spatio-Temporal Confounding

@@ Abstract
In epidemiological studies of air pollution and public health, estimating the health impact of exposure to air pollution may be hindered by the unknown functional form of the exposure–outcome association and by unmeasured confounding factors that are linked to both exposure and outcome. These challenges are especially relevant in spatio-temporal analyses, where their joint exploration remains limited. To study the effects of fine particulate matter on mortality among elderly people in Italy, we propose a Bayesian spatial dynamic generalized linear model that captures the non-linear exposure–outcome association and decomposes the exposure effect across fine and coarse spatio-temporal scales of variation. Together, these features allow reducing the spatio-temporal confounding bias and recovering the shape of the association, as demonstrated through simulation studies. The real-data analysis reveals a clear temporal pattern in the exposure effect, with peaks during summer months. We argue that this finding may be due to interactions of particulate matter with air temperature and unmeasured confounders.
Keywords: Air pollution, Bayesian, Dynamic generalized linear model, Non-linearity, Spatio-temporal confounding.
@@ 1 Introduction
Air pollution is a major environmental exposure that increases the risk for cardiovascular and respiratory disease, among others. There is a very rich literature demonstrating the short- and long-term consequences of air pollution on health outcomes (e.g.,Dominici et al. (2002), Janes et al. (2007), Zeger et al. (2008), Greven et al. (2011), Dai et al. (2014), Di et al. (2017), Wu et al. (2020), Brunekreef et al. (2021b), Renzi et al. (2021), Dominici et al. (2022), Gamerman et al. (2022), Requia et al. (2024)). Collectively, these health effects translate into a sizeable contribution to population-level morbidity and premature mortality. Though the impacts of air pollution on public health are well established, quantifying those impacts on a specific population is challenging. Exposure and risk are inherently spatio-temporal phenomena, where variations in particulate types and levels, the heterogeneity in population characteristics and dynamics, as well as the context in which measurements occur all influence and complicate both estimation and interpretation.
The ultimate goal of this paper is to estimate the small-scale spatio-temporal association between all-cause mortality and air pollution for two Italian regions, Piemonte and Lombardia, over the period from 2018 to 2022 (see Section 2). We limit our study to particulate matter with a diameter less than $2.5\mu m$ (PM2.5), which is known to pose the greatest health risk.
@@
As with any regression modeling approach, we must be concerned about the risk of potential bias arising from model misspecification. In the context of our spatio-temporal regression, the exposure–outcome relationship is challenging to model because both the exposure and the health risk vary across time and space. For example: (i) the physical properties and chemical composition of the PM mixture exhibit substantial variation across locations and over time, influenced by environmental conditions, human activities, and episodic phenomena such as sandstorms and wildfires (Chen et al., 2013, Kelly and Fussell, 2012, Masselot et al., 2022, Peng et al., 2005, Zhou et al., 2021); (ii) human susceptibility to PM2.5 is not static, but is modulated by temporally and spatially varying factors, including extreme temperatures, which can exacerbate health risks in specific regions or seasons (Bergmann et al., 2020, Peng et al., 2005, Roberts, 2004, Li et al., 2017); (iii) the relationship between true and measured exposure varies across time and space, with seasonal factors — such as increased summer ventilation — potentially reducing exposure measurement error (Bergmann et al., 2020, Peng et al., 2006, Roberts, 2004, Stafoggia et al., 2008); (iv) factors such as lockdowns, the use of facial masks, physical distancing and other COVID-19 related measures have significantly altered people’s habits and interactions with the environment since 2020; and (v) morphological and climatic differences across health districts may influence the exposure–outcome relationship.
Often the simplifying assumption of a constant and linear exposure–outcome relationship is made. However, if interactions with confounding variables are ignored, this simple relationship may be inadequate. We acknowledge that all models are inherently flawed, but some of them are so misspecified that they can lead to incorrect inferential statements (Gelman and Shalizi, 2013).
There are several established strategies for incorporating a non-linear exposure–outcome relationship. One option is to adopt a functional approach and estimate an exposure–response function (e.g., Brunekreef et al., 2021b, Cork et al., 2023, Dominici et al., 2022, Renzi et al., 2021). However, the inherent spatial and temporal variation in the relationship between PM2.5 and mortality is typically not accounted for. Alternatively, and the strategy used in this work, one can specify a model that explicitly incorporates spatial and temporal variation in the relationship of interest (Gelfand et al., 2003).
Within the second strategy, several modeling approaches have been proposed; however, they present limitations within our framework. For example, the exposure effect can be modeled either as a periodic function (Chen et al., 2013, Peng et al., 2005) or as a season-specific step function (Dai et al., 2014, Klompmaker et al., 2023). However, because the shape of the effect is assumed to be the same across years, these models cannot account for exceptional events such as the outbreak of the SARS-CoV-2 pandemic, which occurred during the period of interest in our study. Further, the abrupt changes in the exposure effect between adjacent seasons implicit in the approaches proposed by Dai et al. (2014) and Klompmaker et al. (2023) are unlikely in practice. The exposure effect depends on meteorological conditions and human activity patterns, which tend to change smoothly over time (Peng et al., 2005, Bhaskaran et al., 2013). Chiogna and Gaetan (2002) and Lee and Shaddick (2008) employ dynamic generalized linear models (DGLMs; West and Harrison, 1997) to analyze the relationship between air pollution and mortality in time series studies; however, their methods are not suitable for spatio-temporal data. Spatio-temporally varying coefficients models (STVCMs; Banerjee et al., 2014, Sahu, 2022) have also been applied in a variety of fields, including healthcare (Cai et al., 2013, Song et al., 2020), as well as air pollution and climate studies (Paez et al., 2008, Bakar et al., 2015). However, these models do not incorporate space-time interaction terms, thus limiting their ability to capture complex spatio-temporal patterns.
In addition to model misspecification bias, we also face potential bias from confounding. When data are spatially correlated, the challenge has received considerable attention in recent years — e.g., Reich et al. (2006), Paciorek (2010), Thaden and Kneib (2018), Papadogeorgou et al. (2019), Keller and Szpiro (2020), Giffin et al. (2021), Gilbert et al. (2021), Nobre et al. (2021), Bobb et al. (2022), Dupont et al. (2022), Khan and Calder (2022), Marques et al. (2022), Zimmerman and Ver Hoef (2022), Azevedo et al. (2023), Guan et al. (2023), Woodward et al. (2024), Gilbert et al. (2025), Wiecha et al. (2025), Wu and Banerjee (2025) and Zaccardi et al. (2025). Gilbert et al. (2021) and Khan and Berrett (2023) note that the challenge of “spatial confounding” is further complicated by the lack of a single, common definition, and summarized several phenomena related to spatial confounding found in the literature.
Though the phenomena summarized in Gilbert et al. (2021) and Khan and Berrett (2023) can be naturally extended to data that are structured in both space and time, the problem of spatio-temporal confounding has received comparatively less attention. Reich et al. (2021) discuss extensions of causal inference methods to settings where exposures and outcomes vary across space and time. Prates et al. (2022) and Adin et al. (2023) propose extensions of methods originally developed for purely spatial settings, inspired by the restricted regression approach (Reich et al., 2006, Hodges and Reich, 2010). Their proposed methods adopt the “random effect collinearity” perspective, where confounding arises from collinearity between exposure and structured random effects (Gilbert et al., 2021).
In this work, we adopt an alternative definition, that is, spatio-temporal confounding is taken as “omitted confounder bias,” which refers to the existence of unmeasured confounding factors that have a spatio-temporal structure affecting both exposure and outcome (Gilbert et al., 2021). In epidemiological studies on the health effects of air pollution, information on some confounders — such as meteorological variables — is typically available. However, others may be unmeasured or difficult to quantify, such as influenza epidemics or changes in population structure (Peng et al., 2006). Failing to fully account for unmeasured confounding can lead to an incorrect inference about the exposure–outcome relationship (due to confounding bias).
We seek a comprehensive data-driven approach that is able to 1) adequately control for spatio-temporal confounding; 2) capture the pattern of the exposure effect in space and/or time; and 3) break down the spacetime-varying coefficients into simpler parts to enhance interpretability. We, therefore, propose the spatial DGLM for confounding (SDGLMC). To implement this structure, we adopt a Bayesian DGLM framework introducing spacetime-varying intercept and slope coefficients into the model specification to account for unmeasured spatio-temporal confounders and non-linear exposure–outcome associations, respectively. Because DGLMs are only locally linear, they can approximate any smooth function of time (Gamerman and Lopes, 2006, Granger, 2008). Further, because DGLMs capture gradual changes and emphasize both seasonal and non-systematic variations (Lütkepohl, 2005, Petris et al., 2009, Bitto and Frühwirth-Schnatter, 2019), SDGLMC can accommodate exceptional events and avoids abrupt shifts in the exposure effect. Following Greven et al. (2011) and Janes et al. (2007), we decompose the exposure effect into two orthogonal components, each of which describes the (non-linear) association with the health outcome at a different scale of variation (see Section 3). The focus of inference is the component that captures the association at a finer scale — where confounding by unmeasured factors is less likely (Greven et al., 2011) — and is assumed to exhibit spatio-temporal dynamics. Furthermore, SDGLMC enables explicit interpretation of each component of these dynamics (i.e., baseline, temporal, spatial, and interaction effects) through an analysis-of-variance-like decomposition (see Section 3.1).
@@
This paper is organized as follows. Section 2 offers a brief description of the data that motivated this study. Sections 3 and 4 provide detailed information about the proposed model and the inferential procedure, respectively. Section 5 reviews existing approaches from the literature, which serve as benchmarks for comparison with the results of the proposed model. In Section 6, simulation studies are conducted to evaluate the ability of SDGLMC to control for spatio-temporal confounding and recover the true association. The results of the real-data application are presented in Section 7. Finally, Section 8 provides insights from the analysis and offers a concluding discussion.
2 Motivation
@@
We analyze the association between daily PM2.5 concentrations and all-cause mortality counts among people aged 65 and older in the Italian regions of Piemonte and Lombardia between 2018 and 2022.
@@
The data were retrieved from publicly available sources; the data sources and processing procedures are detailed in Section A of the Supplementary Material. Mortality counts were collected from the Italian National Institute of Statistics (ISTAT, 2023a) and are observed on an irregular lattice comprising N=117{N}=117 non-overlapping areal units, termed health districts (HDs). We denote the number of deaths (by any cause) that occurred in the $i$-th HD, $i=1,\dots,{N}$, on day $t=1,\dots,{T}$ as $Y_{it}$, where T{T} is the study period (T=1826{T}=1826 consecutive days). PM2.5 concentrations are used as the measure of exposure in HDi on day $t$, denoted as $X_{it}$, and were retrieved from the Copernicus Atmosphere Monitoring Service’s (CAMS) Atmosphere Data Store (INERIS et al., 2022, CAMS, 2023). Temperature and relative humidity measurements were obtained from the ERA5 reanalyses dataset available at the Copernicus Climate Change Service’s (C3S) Climate Data Store (C3S, 2023, Hersbach et al., 2018) and were spatially and temporally aggregated to match the resolution of the health data.
@@
A measure commonly employed for exploratory analyses is the standardized mortality ratio (SMR), which is calculated as $SMR_{it}=Y_{it}/E_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$. The denominator, $E_{it}$, denotes the expected number of deaths and is computed allowing for the underlying age–sex stratification of the population within each HD (see Section A of the Supplementary Material). The daily SMR for each HD averaged across the study period is shown in the left panel of Figure 1, where noticeable spatial clustering is seen. The mean SMR across the study domain is $1.25$. The right panel of Figure 1 shows the long-term average PM2.5 concentrations for each HD, indicating that Lombardia is generally more polluted than Piemonte ($19.41\,\mu g/m^{3}$ and $15.18\,\mu g/m^{3}$, respectively). Additional descriptive statistics are provided in Section B of the Supplementary Material.
@@
Figure 1: Long-term averages of SMR (left) and PM2.5 concentrations (right) across HDs from 2018 to 2022. The thicker solid line separates Piemonte (the westernmost region) from Lombardia (the easternmost region).
@@
Figure 2 shows the observed daily average SMR on the log scale (orange) and the daily average PM2.5 concentration (black) for Lombardia and Piemonte. Both regions exhibit similar seasonal patterns, with the lowest average SMR and PM2.5values in summer — except for the year 2020. Two distinct SMR peaks are observed in 2020, one in April and another in November, corresponding to the first two waves of the SARS-CoV-2 pandemic. The first wave was more severe in Lombardia, while the second one was more pronounced in Piemonte. Notably, these peaks in the SMR time series are not reflected in the PM2.5 time series. In fact, as shown in Table B.3 of the Supplementary Material, the lowest overall annual mean PM2.5 during the study period was observed in 2020.
@@
Figure 2: Time series of log-transformed daily SMR (orange) and PM2.5 concentrations (black) for Piemonte (top) and Lombardia (bottom), spanning 2018–2022. Vertical dashed lines mark calendar years. orange bands indicate national lockdown periods.
@@
Additional features can be seen in Figure 2. For example, the highest SMR values in 2018 occurred in early January, suggesting that the seasonal influenza outbreak reached its peak before the end of 2017. In contrast, the SMR peaked again in February of the following year, suggesting a longer or later influenza outbreak. The increase in mortality observed during summer 2022 may be due to the occurrence of prolonged heat waves — the average daily temperature time-series has its absolute maximum in July 2022 — or an uncommon summer influenza outbreak.
Figure 2 illustrates the complexity of SMR dynamics over time. Some features clearly reflect factors unrelated to PM2.5. The departures highlight the need for models that can account for the effects of measured and unmeasured confounders. Failing to do so may lead to incorrect inferences. The modeling approach proposed in the next section is designed to address this issue.